OpenStudy (anonymous):

You are considering buying one of two brands of barbecue grills. Brand F costs $650 and will last for about fifteen years. Brand G costs $200 and will last for about five years, so you will need to buy three of them over the years to equal one Brand F grill. In either case, you plan to pay for the grill with your credit card, which has an interest rate of 13.01%, compounded monthly. You will pay off a Brand F grill in five years of monthly payments, and you will pay off a Brand G grill in three years of monthly payments.

OpenStudy (anonymous):

Assuming that you have no other purchases on your credit card, over a fifteen-year period, which kind of grill will be cheaper, and how much cheaper will it be? (Round all dollar values to the nearest cent.) a. Brand G will end up costing $159.48 less than Brand F. b. Brand G will end up costing $50.00 less than Brand F. c. Brand F will end up costing $134.28 less than Brand G. d. Brand F will end up costing $91.08 less than Brand G.

OpenStudy (anonymous):

Can someone please help, I don't really have a teacher and the 10 min video didn't explain how to go about doing this. If someone could explain it id appreciate it.

OpenStudy (anonymous):

M = P * ( J / (1 - (1 + J)^ -N)). M: monthly payment P: principal or amount of loan J: monthly interest; annual interest divided by 100, then divided by 12. N: number of months of amortization, determined by length in years of loan. Applying this to Brand F. M = 650 * (0.01084166666 / (1 - (1 + 0.01084166666)^-(5 * 12))) M = 14.79 And over the 5 years that is 14.79 * 5 * 12 = $887.4 Applying this to Brand G M = 200 * (0.01084166666 / (1 - (1 + 0.01084166666)^-(3 * 12))) M = 6.73 And over the 3 years that is 6.73 * 3 * 12 = 242.28 Doing this three times gives a total price of 726.84 So Brand G is cheaper by 161.16

OpenStudy (anonymous):

i am just trying this answer i got from some site i thought it is urgent so i posted here

OpenStudy (anonymous):

thank you, i appreciate it :)

OpenStudy (anonymous):

welcome :)